# Hamiltonian reductions as affine closures of cotangent bundles

### v1

@@ 5. Differential operators on orbit spaces
An element $\phi\in S^{k}(Z):=H^{0}(Z_{\operatorname{reg}},{\sf S}^{k}T_{Z_{\operatorname{reg}}})$ is called an order $k$ symbol. The ring $\mathcal{D}(Z)$ has a filtration $\{\mathcal{D}^{k}(Z)\}$ by order of differentiations, and the associated graded ring $\operatorname{gr}\mathcal{D}(Z)$ is commutative. Then the order $k$ symbol map gives a linear map $\sigma_{Z,k}\colon\mathcal{D}^{k}(Z)\rightarrow S^{k}(Z)$, which defines an injective homomorphism of graded algebras, called the graded symbol map:
@@
We will consider the case where $Z$ is a quotient, i.e. Z=Y//GZ=Y/\!/G, where $Y$ is an irreducible non-singular affine $G$-variety. This problem was investigated in [Sch95]. Recall that there exists a canonical morphism $(\pi_{Y})_{*}\colon\mathcal{D}(Y)^{G}\rightarrow\mathcal{D}(Z)$ preserving the filtrations by order, where $\pi_{Y}\colon Y\rightarrow Z$ is the canonical quotient morphism. Then $(\pi_{Y})_{*}$ is called graded surjective if the induced graded morphism $\operatorname{gr}(\pi_{Y})_{*}\colon\operatorname{gr}(\mathcal{D}(Y)^{G})\longrightarrow\operatorname{gr}(\mathcal{D}(Z))$ is surjective.
@@ A.2. Auxiliary constructions
Let $S_{y}(\cong S)$ be the fibre of $Y=G\times_{H}S\rightarrow G/H$ over $[H]$. Then $y\in S_{y}$. Let $\mathfrak{m}_{S_{y},y}\subset\mathcal{O}(S_{y})$ be the maximal ideal corresponding to $y$. Then we have a composition of linear map
@@
Let $L$ be the subspace of $\mathfrak{m}_{S_{y},y}\subset\mathcal{O}(S_{y})$ generated by $H_{\xi}$ with $\xi\in\mathfrak{l}$. By Lemma A.2, we have $L\cap\mathfrak{m}_{S_{y},y}^{2}=\{0\}$. In particular, since $L$ and $\mathfrak{m}^{2}_{S_{y},y}$ are $H$-stable subspaces of $\mathfrak{m}_{S_{y},y}$ and $H$ is reductive, we may choose an $H$-stable complement $F$ of $L\oplus\mathfrak{m}^{2}_{S_{y}}$ in $\mathfrak{m}_{S_{y},y}$. Then the natural inclusion
@@
induces a dominant morphism $\nu\colon S_{y}\rightarrow(F\oplus L)^{*}=F^{*}\oplus L^{*}$ and an linear isomorphism
@@
Clearly $\nu$ is étale at $y$ and it maps the fixed point $y$ to $0\in F^{*}\oplus L^{*}$. Then Luna’s fundamental lemma [Lun75] implies that $\nu$ is an excellent morphism in an open affine $H$-saturated neighborhood of $y$. Thus we may assume that $\nu$ itself is excellent after shrinking $S_{y}$. Now let $U$ be the preimage $\nu^{-1}(F^{*})$. Then $U$ is non-singular since $\nu$ is étale.
@@
Just note that $L$ is generated by the functions $H_{\xi}|_{S_{y}}$, $\xi\in\mathfrak{l}$. ∎
@@
The restriction of $\omega_{Y}$ to $U$ is symplectic in an open neighbourhood of $y$.
@@
Consider the linear map $\alpha|_{\mathfrak{l}}\colon\mathfrak{l}\rightarrow T_{y}Y\rightarrow T_{y}^{*}Y$. Then the image of $\mathfrak{l}$ in $T_{y}Y$ is exactly $T_{y}O_{y}$. Let us identify $\mathfrak{l}$ to its image in $T_{y}^{*}Y$; that is, the subspace generated by $dH_{\xi}(y)$ for $\xi\in\mathfrak{l}$. Let $\operatorname{Ann}(\mathfrak{l})\subset T_{y}Y$ be the annilhilator of $\mathfrak{l}$, which contains the tangent space $T_{y}U$ by the construction of $U$. In particular, since $T_{y}O_{y}$ is isotropic with respect to $\omega_{Y,y}$, a dimension counting then yields

### v2

@@ 5. Differential operators on orbit spaces
An element $\phi\in S^{k}(Z):=H^{0}(Z_{\operatorname{reg}},{\sf S}^{k}T_{Z_{\operatorname{reg}}})$ is called an order $k$ symbol. The ring $\mathcal{D}(Z)$ has a filtration $\{\mathcal{D}^{k}(Z)\}$ according to the order of differentiations, and the associated graded ring $\operatorname{gr}\mathcal{D}(Z)$ is commutative. Then the order $k$ symbol map gives a linear map $\sigma_{Z,k}\colon\mathcal{D}^{k}(Z)\rightarrow S^{k}(Z)$, which defines an injective homomorphism of graded algebras, called the graded symbol map:
@@
We will consider the case where $Z$ is a quotient, i.e. Z=Y//GZ=Y/\!/G, where $Y$ is an irreducible non-singular affine $G$-variety. This problem was investigated in [Sch95]. Recall that there exists a canonical morphism $(\pi_{Y})_{*}\colon\mathcal{D}(Y)^{G}\rightarrow\mathcal{D}(Z)$ that preserves the filtrations by order, where $\pi_{Y}\colon Y\rightarrow Z$ is the canonical quotient morphism. Then $(\pi_{Y})_{*}$ is called graded surjective if the induced graded morphism $\operatorname{gr}(\pi_{Y})_{*}\colon\operatorname{gr}(\mathcal{D}(Y)^{G})\longrightarrow\operatorname{gr}(\mathcal{D}(Z))$ is surjective.
@@ A.2. Auxiliary constructions
Let $S_{y}(\cong S)$ be the fiber of $Y=G\times_{H}S\rightarrow G/H$ over $[H]$. Then $y\in S_{y}$. Let $\mathfrak{m}_{S_{y},y}\subset\mathcal{O}(S_{y})$ be the maximal ideal corresponding to $y$. Then we have a composition of linear maps
@@
Let $L$ be the subspace of $\mathfrak{m}_{S_{y},y}\subset\mathcal{O}(S_{y})$ generated by $H_{\xi}$ with $\xi\in\mathfrak{l}$. By Lemma A.2, we have $L\cap\mathfrak{m}_{S_{y},y}^{2}=\{0\}$. In particular, since $L$ and $\mathfrak{m}^{2}_{S_{y},y}$ are $H$-stable subspaces of $\mathfrak{m}_{S_{y},y}$ and $H$ is reductive, we may choose a $H$-stable complement $F$ of $L\oplus\mathfrak{m}^{2}_{S_{y}}$ in $\mathfrak{m}_{S_{y},y}$. Then the natural inclusion
@@
induces a dominant morphism $\nu\colon S_{y}\rightarrow(F\oplus L)^{*}=F^{*}\oplus L^{*}$ and a linear isomorphism
@@
Clearly $\nu$ is étale at $y$ and maps the fixed point $y$ to $0\in F^{*}\oplus L^{*}$. Then Luna’s fundamental lemma [Lun75] implies that $\nu$ is an excellent morphism in an open affine $H$-saturated neighborhood of $y$. Thus, we may assume that $\nu$ itself is excellent after shrinking $S_{y}$. Now, let $U$ be the preimage $\nu^{-1}(F^{*})$. Then $U$ is non-singular since $\nu$ is étale.
@@
Note that $L$ is generated by the functions $H_{\xi}|_{S_{y}}$, $\xi\in\mathfrak{l}$. ∎
@@
The restriction of $\omega_{Y}$ to $U$ is symplectic in an open neighborhood of $y$.
@@
Consider the linear map $\alpha|_{\mathfrak{l}}\colon\mathfrak{l}\rightarrow T_{y}Y\rightarrow T_{y}^{*}Y$. Then the image of $\mathfrak{l}$ in $T_{y}Y$ is exactly $T_{y}O_{y}$. Let us identify $\mathfrak{l}$ with its image in $T_{y}^{*}Y$; that is, the subspace generated by $dH_{\xi}(y)$ for $\xi\in\mathfrak{l}$. Let $\operatorname{Ann}(\mathfrak{l})\subset T_{y}Y$ be the annilhilator of $\mathfrak{l}$, which contains the tangent space $T_{y}U$ by the construction of $U$. In particular, since $T_{y}O_{y}$ is isotropic with respect to $\omega_{Y,y}$, a dimension count then yields